OpenStudy (anonymous):

How many solutions does the equation 5x + 3x –4 = 10 have? (4 points) Infinitely many Zero Two One

OpenStudy (idealist10):

5x+3x-4=10 8x-4=10 8x=10+4 8x=14 x=14/8 simplify x=7/4 Answer: 7/4, just one solution.

OpenStudy (anonymous):

Thats what I thought thanks

OpenStudy (anonymous):

What is the value of x in the equation 3(2x + 4) = –6? (4 points) 19 1 –3 12 @Idealist10

OpenStudy (idealist10):

3(2x+4)=-6 6x+12=-6 6x=-6-12 6x=-18 x=-18/6 x=-3 Answer: -3

OpenStudy (anonymous):

What is the value of x in the equation 2(2x –5 –4) = 160 –17? (4 points) 40.25 143.0 62.25 80.25

OpenStudy (anonymous):

@Idealist10

OpenStudy (idealist10):

2(2x-5-4)=160-17 2(2x-9)=143 4x-18=143 4x=143+18 4x=

OpenStudy (idealist10):

4x=161 x=161/4 x=40.25 Answer: 40.25

OpenStudy (anonymous):

What is the solution to the following equation? 4(3x –6) + 15 = 27 (4 points) 4 2 5 3

OpenStudy (idealist10):

4(3x-6)+15=27 4(3x-6)=27-15 4(3x-6)=12 3x-6=12/4 3x-6=3 3x=3+6 3x=9 x=9/3 x=3 Answer: 3